OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (1, 7), (10, 1) A. –2/3 B. 3/2 C. 2/3 D. –3/2

OpenStudy (anonymous):

@toomuch2handle

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

do you know the formula of slope

OpenStudy (welshfella):

slope = difference in y coordinates / difference in x coordinates

OpenStudy (foreverandalways13):

ok so the slope formula is y2-y1 over x2-x1 so plug the numbers in and you get -2/3

OpenStudy (anonymous):

No, not A.

OpenStudy (anonymous):

Can you explain it

OpenStudy (anonymous):

\[slope= y _{2}-y _{1}/x _{2}-x _{1}\]

OpenStudy (anonymous):

Slope is rise over run. So, y2-y1/x2-x1

OpenStudy (welshfella):

(1 - 7) / (10 - 1) = -6 / 9 = -2/3

OpenStudy (anonymous):

consider first points as x1,y1 and second one as x2, y2

OpenStudy (anonymous):

So, 1-10/7-1. Solve that and you will have the answer.

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

i think if you try than it's better for understanding even i solve all

OpenStudy (anonymous):

No, the answer would be -9/6. Comes to....... now you figure the rest out.

OpenStudy (anonymous):

It would be -3/2, or -1.5. So, D

OpenStudy (welshfella):

No the snwer is A x1 = 1 x2 = 10 y1= 7 y2 = 1 y2 - y1 / x2 - x1 = 1/7 / 10-1 = -6/9 = -2/3